Kayla took a 4-year auto-loan of $20,000. Assume the interest rate is 8 percent per year and the loan agreement calls for fixed principal payment per year. What is the interest payment in year 2?
Answer #1

Fixed Principal repaid Per Year = Auto Loan amount/Number of Years =20000/4 =5000
Principal after year 1 = Loan - Principal Per year = 20000 -5000 =15000
Interest payment in year 2 = Principal in Year 1* Interest rate =15000*8/100 =1200